OpenStudy (anonymous):

Can someone help me prove this:

OpenStudy (anonymous):

\[a ^{x} = e ^{(x)\ln a }\]

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@ganeshie8

OpenStudy (accessdenied):

Did you have any ideas on how to approach this so far?

OpenStudy (anonymous):

I only know the basic properties of natural logs. for example: e^(lna) = a

OpenStudy (accessdenied):

Yup. How about the property for a power to a power? \( (x^a)^b = x^{ab} \)

OpenStudy (anonymous):

Yes. I am also familiar with that property. My question here is how to apply these properties to the formula provided.

OpenStudy (accessdenied):

I think we should try to convert the right-hand side into the left. \( a^x = e^{x \ln a} \) We have x times ln a. So how about we apply the power to a power property and rewrite it so that: \( e^{x \ln a} = e^{(\ln a ) x} = \left( e^{\ln a} \right)^x \)

OpenStudy (anonymous):

@AccessDenied Ohhhh! I knew that but this question dazzled me for while. Thanks a lot mate, I really appreciate your effort.

OpenStudy (accessdenied):

Glad to help! :)